Grade upto college level Physical Chemistry

The composition of a sample of wustite is Fe0.93O1.00. What percentage of the iron is present in the form of Fe (III)?

Amit Saxena
12 Years agoGrade upto college level

2 Answers

Navjyot Kalra
12 Years ago
In pure iron oxide (FeO), iron and oxygen are present in the ratio 1 : 1.
However, here number of Fe2+ present = 0.93
Or No. of Fe2+ ions missing = 0.07
Since each Fe2+ ion has 2 positive charge, the total number of charge due to missing (0.07) Fe2+ ions = 0.07 * 2 = 0.14
To maintain electrical neutrality, 0.14 positive charge is compensated by the presence of Fe3+ ions. Now since, replacement of one Fe2+ ion by one Fe3+ ion increases one positive charge, 0.14 positive charge must be compensated by the presence of 0.14 Fe3+ ions.
In short, 0.93 Fe2+ ions have 0.14 Fe3+ ions
100 Fe2+ ions have = 0.14/0.93 * 100 = 15.05%
Kushagra Madhukar
6 Years ago
Dear student,
Please find the attached solution to your question.
 
let us take x be the number of Fe3+ ions in 100moles of Fe0.93O1
Hence, moles of Fe2+ = 100 x 0.93 – x = 93 – x
Since it is a neutrally charged compound.
The net charge on it will be balanced
Hence, 2(93 – x) + 3x = 2 x 100
186 + x = 200
or, x = 14
Hence percentage of Fe3+ = 14/93 x 100 = 15.05%
 
Hope this helps.
Thanks and regards,
Kushagra